Bob is three times as old as Jenny, who is five times as old as Lou. If Bob is x
years old, then in ten years how old will Lou be in terms of x?
a) x/(5+10)
b) (x+2)/2
c) (x+6)/3
d) (3x)/(5+10)
e) x/(3+10)

Found 3 solutions by josgarithmetic, Edwin McCravy, greenestamps:
Answer by josgarithmetic(39618)

PERSON            AGE
Bob                x
Jenny              x/3
Lou                x/15

You can find the choice that works.

Answer by Edwin McCravy(20056)

This problem is terribly botched!

Bob is three times as old as Jenny, who is five times as old
as Lou.
This means that Bob, who is x, is 15 times as old as Lou.
So if we kept everybody under 100 years old, Lou could only 
be 1,2,3,4,5 or 6 now. which would make him 11,12,13,14,15,
or 16 in 10 years. So here are the only possibilities:

                 Lou
                  in
            Bob   10   choice  choice  choice  choice  choice 
Lou  Jenny   x   years   a)     b)       c)      d)      e)
 1     5    15    11     1      8.5       7      3      1.1  
 2    10    30    12     2     16        12      6      2.3
 3    15    45    13     3      23.5     17      9      2.5
 4    20    60    14     4      31       22     12      4.6
 5    25    75    15     5      38.5     27     15      5.8
 6    30    90    16     6      46       32     18      8.1

If Bob is x=30, the correct choice will be c)
If Bob is x=75, the correct choice will be d)
But for Bob being any other age, all the choices are wrong.

This is a monstrosity of a problem, but I got my kicks by
analyzing it. lol

Edwin

Answer by greenestamps(13200)



The question doesn't ask for any ages; it only asks for an expression for Lou's age in terms of x, which is Bob's age.

x = Bob
(1/3)x = Jenny (Bob is 3 times as old as Jenny)
(1/15)x = Lou (Jenny is 5 times as old as Lou)

(1/15)x+10 = Lou's age in 10 years

None of the given answer choices is equivalent to that, so the problem is faulty.